OpenStudy (anonymous):

Annie has a compost pile that produces 12 pounds of compost a day from yard waste. Which function expresses the amount of compost (c) as a function of the number of days (d)? The function representing the relationship between the amount of compost and the number of days is ____ , with the input as ____.

OpenStudy (anonymous):

The first blank has the options: c=f(d)=12d d=(f)c=12m c=f(d)=12+d The second blank has the options: d c 12